# Resonances and resonance expansions for $\delta$-interactions on the half-space

### v2

@@ 3. Resonances of point interactions in the half-space
This system is overdetermined and so can have a solution only for certain values of $\alpha$ and $y_{3}$. It is easy to see that if $k$ is positive odd or negative even, then no solution exists regardless of $\alpha$ and $y_{3}$ ($a$ would be in one of the intervals mentioned above for which no solution exists); instead, if $k$ is non-negative even and $\alpha=\frac{1}{8\pi y_{3}}\ln{\Bigl{(}\frac{\pi}{2}+k\pi\Bigr{)}}$, then there is a resonance in $z=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}-\frac{i}{2y_{3}}\ln{\Bigl{(}\frac{\pi}{2}+k\pi\Bigr{)}}$; while, if $k$ is negative odd and $\alpha=\frac{1}{8\pi y_{3}}\ln{\Bigl{(}-\frac{\pi}{2}-k\pi\Bigr{)}}$, then there is a resonance in $z=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}-\frac{i}{2y_{3}}\ln{\Bigl{(}-\frac{\pi}{2}-k\pi\Bigr{)}}$. This proves part v) and vi). After treating the exceptional cases, we turn to part $i)$ and $ii)$. As previously, we locate resonances thanks to the equation $\Gamma_{\alpha,y}^{\textup{D}}(z)=0$. This leads to Eq. 3.3. We divide side by side the second equation for the first one. We notice that all the values leading to a vanishing denominator have been considered in Proposition 3.1 and hence performing this ratio does not exclude any resonance not already considered. Doing so, we get $\tan{(2y_{3}a)}=-\frac{a}{4\pi\alpha+b}$, which, solved for $b$ gives
@@
Substituting $b$ in the second equation of Eq. 3.3, we get the following equation in $a$
(3.5) $g(a)=\frac{\sin{(2y_{3}a)}}{2y_{3}a}=e^{-8\pi\alpha y_{3}-\frac{2y_{3}a}{\tan{(2y_{3}a)}}}=h(a).$
Now we claim that in each interval of the form $\Bigl{(}\frac{k\pi}{y_{3}},\frac{(2k+1)\pi}{2y_{3}}\Bigr{)}$ with $k=1,2,\dots$, Eq. 3.5 has exactly one solution.
Let $a_{k}$ be the smallest solution in the $k-th$ interval. We distinguish the two cases $a_{k}\geq M_{k}$ and $a_{k}<M_{k}$, where $M_{k}$ is the maximum of $g$ in $\Bigl{(}\frac{k\pi}{y_{3}},\frac{(2k+1)\pi}{2y_{3}}\Bigr{)}$. If $a_{k}\geq M_{k}$, the solution is unique because in $\Bigl{[}a_{k},\frac{(2k+1)\pi}{2y_{3}}\Bigr{)}$, $h$ is known to be increasing and $g$ is decreasing thanks to Lemma 3.1 . If $a_{k}<M_{k}$, given that $M_{k}\in\Biggl{(}\frac{k\pi}{y_{3}},\frac{\frac{\pi}{2}+2k\pi}{2y_{3}}\Biggr{)}$, the function $r(a)=h(a)-g(a)$ is strictly convex in $\Biggl{(}\frac{k\pi}{y_{3}},M_{k}\Biggr{)}$ by the just above mentioned lemmata. But this, together with $\lim_{a\to\frac{k\pi^{+}}{y_{3}}}r(a)=0$ implies that at there is at most a zero for $r$ in $\Biggl{(}\frac{k\pi}{y_{3}},M_{k}\Biggr{)}$. In fact, if there were two internal zeros $a_{k}<\tilde{a}_{k}$, there would be two numbers $b_{k}$ and $\tilde{b}_{k}$ such that $\frac{k\pi}{y_{3}}<b_{k}<a_{k}<\tilde{b}_{k}<\tilde{a}_{k}<M_{k}$ and $r^{\prime}(b_{k})=r^{\prime}(\tilde{b}_{k})=0$. This contradicts the function $r$ being convex ($r^{\prime}$ is strictly increasing), there is a solution at most. For each interval, this possible unique solution in fact exists. That’s because the conditions $h\Bigl{(}\frac{k\pi^{+}}{y_{3}}\Bigr{)}=g\Bigl{(}\frac{k\pi}{y_{3}}\Bigr{)}=0$, $h^{\prime}\Bigl{(}\frac{k\pi^{+}}{y_{3}}\Bigr{)}=0<g^{\prime}\Bigl{(}\frac{k\pi}{y_{3}}\Bigr{)}=\frac{y_{3}}{k\pi}$, $h\Bigl{(}\frac{\pi^{-}+2k\pi}{2y_{3}}\Bigr{)}=+\infty$ and $g\Bigl{(}\frac{\pi^{-}+2k\pi}{2y_{3}}\Bigr{)}=0$ ensure that $r$ changes sign throughout the interval.
It remains to consider the interval $\Bigl{[}0,\frac{\pi}{2y_{3}}\Bigr{)}$. Here $g$ is monotone decreasing and $h$ is monotone increasing, so, at most there is one solution. The solution actually exists only if $\lim_{a\to 0}r(a)<0$, which happens if and only if $\alpha>-\frac{1}{8\pi y_{3}}$. This proves part $ii)$.
For each solution of Eq. 3.5, we can recover the corresponding imaginary part through the relation $b=\frac{1}{2y_{3}}\ln{g(a)}$, which corresponds to formula (3.1). This finally proves part $i)$ of the Proposition. ∎
@@
It is well known [31] that for Schrödinger operators $-\Delta+V$, zero energy resonances and zero energy eigenvalues can be distinguished by means of the different low energy behavior of the resolvent. Namely, the resolvent $(-\Delta+V-z^{2})^{-1}$ behaves as $\frac{c_{-2}}{z^{2}}$ at a zero energy eigenvalue and as $\frac{c_{-1}}{z}$ at a zero energy resonance, for certain bounded operators $c_{-1},c_{-2}$, bounded in suitable operator topologies. The same is true in the present example, not covered by the standard theory. Let us consider the Dirichlet case. At the critical value $\alpha_{\textup{c}}=-\frac{1}{8\pi y_{3}}$ one has
@@
Concerning the multiplicity of the resonances, we notice the following. All the resonance values considered as roots of the "characteristic system" (3.3) (and the analog Neumann case) are simple, but the zero resonance in the Dirichlet case, where the root is double. In the latter case, one has an eigenvalue with the corresponding eigenfunction (3.4), and no other solution. This because the corresponding boundary value problem implies that the the regular part of the resonance function satisfies the Laplace equation on the half-plane, which enjoys uniqueness thanks to the Maximum principle (see for example Lemma 2.1 in [9]).
Remark 3.8.
[Bifurcation of eigenvalues and resonances at zero energy] Two different pictures show up in the Dirichlet and Neumann case as regards the bifurcation of eigenvalues and resonances at the origin of the complex plane. In the Neumann case, things are simple: a resonance (an anti-bound state more precisely) moves on the negative imaginary axis when $\alpha\in(\frac{1}{8\pi y_{3}},\infty)$, and disappears at zero. In the Dirichlet case, a couple bound state- resonance separate at $z=0$ moving respectively on the positive imaginary axis (eigenvalue) and on the negative imaginary axis (anti-bound state); the couple bifurcates from the branch of resonances colliding at zero and described in Proposition 3.1, iv).
Remark 3.9.
As noticed in Remark 2.5, operators $-\Delta_{\alpha,y}^{\textup{D}}$ and $-\Delta_{\alpha,y}^{\textup{N}}$ cannot be obtained from a combination of point interactions in $\mathbb{R}^{3}$. However, they could have in principle the same point spectrum or set of resonances of a combination of many center point interactions. For symmetry reasons the only possible candidate is a 2-center point interaction. Choosing the distance of the two centers as $2y_{3}$ and the same $\alpha$ one shows that the set of spectrum and resonances of this last model equals the union of the set of zeroes of $\Gamma_{\alpha,y}^{\textup{D}}(z)$ and of $\Gamma_{\alpha,y}^{\textup{N}}(z)$ (see [2] Chapter II.1, Theorem 1.1.4, in particular formula 1.1.53, giving the set of eigenvalues and resonances for N-center point interactions). Other combinations of distances and $\alpha$’s do not give point spectra comparable with the ones here obtained for the half-space.
@@ 4. Asymptotics of the Resonance Counting Function
In the present Section we study the asymptotic behaviour of the resonance counting function for the one center Laplacian in the half-space. Consider the set $\Sigma\Bigl{(}-\Delta_{\alpha,y}^{\textup{D}}\Bigr{)}$ of complex $z$ such that $z^{2}$ is an eigenvalue or a resonance of $-\Delta_{\alpha,y}^{\textup{D}}$. Let’s define
$\mathcal{N}_{-\Delta_{\alpha,y}^{\textup{D}}}(R)=\#\Set{z\in\Sigma\Bigl{(}-\Delta_{\alpha,y}^{\textup{D}}\Bigr{)}}{\left\lvert z\right\rvert<R}$
the counting function for the the elements of $\Sigma\Bigl{(}-\Delta_{\alpha,y}^{\textup{D}}\Bigr{)}$ counted with the appropriate multiplicity. We now prove that (see [34, 7] for an analogous result for $N$-point interactions in $\mathbb{R}^{n}$)
@@
$\forall\alpha\in\mathbb{R}$ and $y\in\mathbb{R}^{3}_{+}$ it holds that
(4.1) $\mathcal{N}_{-\Delta_{\alpha,y}^{\textup{D}}}(R)\sim 2\frac{y_{3}}{\pi}R\ \ \text{as}\ \ \ R\to\infty\ .$
@@
Let $z_{k}^{2}$ be the resonance of $-\Delta_{\alpha,y}^{\textup{D}}$ such that $\operatorname{Re}z_{k}:=a_{k}\in\Bigl{(}\frac{k\pi}{y_{3}},\frac{(2k+1)\pi}{2y_{3}}\Bigr{)}$ and $\{a_{k}\}_{k=1,2,\dots}$ the associated sequence. We start by proving that
(4.2) $a_{k}\sim\frac{(2k+1/2)\pi}{2y_{3}}\ \ \ \ \text{as}\ \ \ k\to+\infty.$
We observe that, even if it can occur that $a_{1}\in\Bigl{(}\frac{5\pi}{4y_{3}},\frac{3\pi}{2y_{3}}\Bigr{)}$, it still happens that, definitely in $k$, $a_{k}\in\Bigl{(}\frac{k\pi}{y_{3}},\frac{(2k+1/2)\pi}{2y_{3}}\Bigr{)}$. This is true because, recalling Eq. 3.5, we have that
@@
and so there exists a $\overline{k}$, such that $\forall k\geq\overline{k}$, $g<h$ when both evaluated at the middle point of the $k$-th interval, which implies that the unique solution is in the first half of the interval.
@@
Both sequences are vanishing, but $\{h(w_{l})\}$ faster than $\{g(w_{l})\}$, which implies that for sufficiently large $l$, $g(w_{l})>h(w_{l})$ and so the solution in the $(\overline{k}+l)$-th interval must be in $\Bigl{(}w_{l},\frac{(2(\overline{k}+l)+1/2)\pi}{2y_{3}}\Bigr{)}$. The arbitrarity of $w$ in $\Bigl{(}\frac{\overline{k}\pi}{y_{3}},\frac{(2\overline{k}+1/2)\pi}{2y_{3}}\Bigr{)}$ leads to Eq. 4.2.
So we have that $a_{k}=\frac{(2k+1/2)\pi}{2y_{3}}+o(1)$ for $k\to+\infty$. We use this property to estimate $\mathcal{N}_{-\Delta_{\alpha,y}^{\textup{D}}}(R)$ as $R\to+\infty$ excluding at most finitely many resonances. This is equivalent to find the highest $k$ for which
@@
(4.3) $s=\sqrt{4y_{3}^{2}R^{2}-\ln^{2}{(s)}}+o(1).$
We see that $s<2y_{3}R$ and $s=O(R)$. This implies that $\ln^{2}{(s)}=O\Bigl{(}\ln^{2}(R)\Bigr{)}$. Now we substitute this in Eq. 4.3, obtaining
@@
[…]
@@
(4.4) $\mathcal{N}_{-\Delta_{\alpha,y}^{\textup{D}}}(R)\sim 2\Biggl{\lfloor}\frac{y_{3}}{\pi}R-\frac{1}{4}\Biggr{\rfloor}$
giving finally equation (4.1) in the statement ∎
@@
The analogous result proved in [16] for Schrödinger operators $-\Delta+V$ with bounded potentials with compact support in $\mathbb{R}^{n}$ gives
$\lim_{R\to+\infty}\frac{\ln{\Bigl{(}\mathcal{N}_{-\Delta+V}(R)\Bigr{)}}}{\ln{R}}=n$
showing a dependence on the dimension different from the one obtained here. This different behavior was already noticed in [34] for finitely many point interactions in $\mathbb{R}^{3}$, and compared with the analogous behavior on quantum graphs. See also the earlier work [43]. The reason of this anomalous behavior is not immediately interpretable, in view of the known fact that point interactions are singular limit of standard Schrödinger operators with rescaled smooth potentials (see for example [2]).

### v3

@@ 3. Resonances of point interactions in the half-space
This system is overdetermined and so can have a solution only for certain values of $\alpha$ and $y_{3}$. It is easy to see that if $k$ is positive odd or negative even, then no solution exists regardless of $\alpha$ and $y_{3}$ ($a$ would be in one of the intervals mentioned above for which no solution exists); instead, if $k$ is non-negative even and $\alpha=\frac{1}{8\pi y_{3}}\ln{\Bigl(\frac{\pi}{2}+k\pi\Bigr)}$, then there is a resonance in $z=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}-\frac{i}{2y_{3}}\ln{\Bigl(\frac{\pi}{2}+k\pi\Bigr)}$; while, if $k$ is negative odd and $\alpha=\frac{1}{8\pi y_{3}}\ln{\Bigl(-\frac{\pi}{2}-k\pi\Bigr)}$, then there is a resonance in $z=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}-\frac{i}{2y_{3}}\ln{\Bigl(-\frac{\pi}{2}-k\pi\Bigr)}$. This proves part v) and vi).
After treating the exceptional cases, we turn to part $i)$ and ii)ii). As previously, we locate resonances thanks to the equation $\Gamma_{\alpha,y}^{\textup{D}}(z)=0$. This leads to Eq.˜3.8. We divide side by side the second equation for the first one. We notice that all the values leading to a vanishing denominator have been considered previously and hence performing this ratio does not cause loss of solutions. So, we get $\tan{(2y_{3}a)}=-\frac{a}{4\pi\alpha+b}$, which, solved for $b$ gives
@@
Substituting $b$ in the second equation of Eq.˜3.8, we get the following equation in $a$
$g(a)=\frac{\sin{(2y_{3}a)}}{2y_{3}a}=e^{-8\pi\alpha y_{3}-\frac{2y_{3}a}{\tan{(2y_{3}a)}}}=h(a).$ (3.10)
Now we claim that in each interval of the form $\Bigl(\frac{k\pi}{y_{3}},\frac{(2k+1)\pi}{2y_{3}}\Bigr)$ with $k=1,2,\dots$, Eq.˜3.10 has exactly one solution.
Let $a_{k}$ be the smallest solution in the kthk-th interval. We distinguish the two cases $a_{k}\geq M_{k}$ and $a_{k}<M_{k}$, where $M_{k}$ is the maximum of $g$ in $\Bigl(\frac{k\pi}{y_{3}},\frac{(2k+1)\pi}{2y_{3}}\Bigr)$. If $a_{k}\geq M_{k}$, the solution is unique because in $\Bigl[a_{k},\frac{(2k+1)\pi}{2y_{3}}\Bigr)$, $h$ is known to be increasing and $g$ is decreasing thanks to Lemma˜3.1 . If $a_{k}<M_{k}$, given that $M_{k}\in\Biggl(\frac{k\pi}{y_{3}},\frac{\frac{\pi}{2}+2k\pi}{2y_{3}}\Biggr)$, the function r(a)=h(a)g(a)r(a)=h(a)-g(a) is strictly convex in $\Biggl(\frac{k\pi}{y_{3}},M_{k}\Biggr)$ by the just above mentioned lemma. But this, together with $\lim_{a\to\frac{k\pi^{+}}{y_{3}}}r(a)=0$ implies that at there is at most a zero for $r$ in $\Biggl(\frac{k\pi}{y_{3}},M_{k}\Biggr)$. In fact, if there were two internal zeros $a_{k}<\tilde{a}_{k}$, there would be two numbers $b_{k}$ and $\tilde{b}_{k}$ such that $\frac{k\pi}{y_{3}}<b_{k}<a_{k}<\tilde{b}_{k}<\tilde{a}_{k}<M_{k}$ and $r^{\prime}(b_{k})=r^{\prime}(\tilde{b}_{k})=0$. This contradicts the function $r$ being convex ($r^{\prime}$ is strictly increasing); then there is a solution at most. For each interval, this possible unique solution in fact exists. This is true because the conditions
$h\Bigl(\frac{k\pi^{+}}{y_{3}}\Bigr)=g\Bigl(\frac{k\pi}{y_{3}}\Bigr)=0,\quad h^{\prime}\Bigl(\frac{k\pi^{+}}{y_{3}}\Bigr)=0<g^{\prime}\Bigl(\frac{k\pi}{y_{3}}\Bigr)=\frac{y_{3}}{k\pi},\quad h\Bigl(\frac{\pi^{-}+2k\pi}{2y_{3}}\Bigr)=+\infty\quad\textup{and}\quad g\Bigl(\frac{\pi^{-}+2k\pi}{2y_{3}}\Bigr)=0$ (3.11)
ensure that $r$ changes sign throughout the interval.
It remains to consider the interval $\Bigl[0,\frac{\pi}{2y_{3}}\Bigr)$. Here $g$ is monotone decreasing and $h$ is monotone increasing, so, at most there is one solution. The solution actually exists only if $\lim_{a\to 0}r(a)<0$, which happens if and only if $\alpha>-\frac{1}{8\pi y_{3}}$. This proves part ii)ii).
For each solution of Eq.˜3.10, we can recover the corresponding imaginary part through the relation $b=\frac{1}{2y_{3}}\ln{g(a)}$, which corresponds to formula (3.4). This finally proves part $i)$ of the proposition. ∎
@@
It is well known (see [38]) that for Schrödinger operators $\Delta+V$, zero energy resonances and zero energy eigenvalues can be distinguished by means of the different low energy behavior of the resolvent. Namely, the resolvent $(-\Delta+V-z^{2})^{-1}$ behaves as $\frac{c_{-2}}{z^{2}}$ at a zero energy eigenvalue and as $\frac{c_{-1}}{z}$ at a zero energy resonance, for certain operators $c_{-1},c_{-2}$ that are bounded in suitable operator topologies. The same is true in the present example, not covered by the standard theory. Let us consider the Dirichlet case. At the critical value $\alpha_{\textup{c}}=-\frac{1}{8\pi y_{3}}$ one has
@@
[Bifurcation of eigenvalues and resonances at zero energy] Two different pictures show up in the Dirichlet and Neumann case as regards the bifurcation of eigenvalues and resonances at the origin of the complex plane. In the Neumann case, things are simple: a resonance (an anti-bound state more precisely) moves on the negative imaginary axis when $\alpha\in(\frac{1}{8\pi y_{3}},\infty)$, and becomes an eigenvalue after crossing the threshold $\alpha=\frac{1}{8\pi y_{3}}$. At threshold, a zero energy resonance appears. In the Dirichlet case, a couple bound state-resonance separate at $z=0$ moving respectively on the positive imaginary axis (eigenvalue) and on the negative imaginary axis (anti-bound state); the couple bifurcates from the branch of resonances colliding at zero, which is an eigenvalue, as described in Proposition 3.4, iv). In both cases, with a different phenomenology, it should be noted that the eigenvalue persists (see for this terminology the recent paper [18]) and it does not disappear. A general understanding of the phenomenon of persistence/disappearing of eigenvalues at threshold seems still lacking.
@@ 4. Asymptotics of the Resonance Counting Function
In the present section we study the asymptotic behaviour of the resonance counting function for the one center Laplacian in the half-space. Consider the set $\Sigma\Bigl(-\Delta_{\alpha,y}^{\textup{D}}\Bigr)$ of resonances of $\Delta_{\alpha,y}^{\textup{D}}$. Let’s define
[…]
the counting function for the the elements of $\Sigma\Bigl(-\Delta_{\alpha,y}^{\textup{D}}\Bigr)$, counted with the appropriate multiplicity, contained in the disk of radius $R$ centered in zero. We now prove that (see [40, 7] for an analogous result for $N$-point interactions in $\mathbb{R}^{n}$)
@@
For all $\alpha\in\mathbb{R}$ and $y\in\mathbb{R}^{3}_{+}$ it holds that
$\lim_{R\to+\infty}\frac{\mathcal{N}_{-\Delta_{\alpha,y}^{\textup{D}}}(R)}{R}=2\frac{y_{3}}{\pi}.$ (4.1)
@@
Let $z_{k}$ be the resonance of $\Delta_{\alpha,y}^{\textup{D}}$ such that $\operatorname{Re}z_{k}:=a_{k}\in\Bigl(\frac{k\pi}{y_{3}},\frac{(2k+1)\pi}{2y_{3}}\Bigr)$ and $\{a_{k}\}_{k=1,2,\dots}$ the associated sequence. We start by proving that
$\lim_{k\to+\infty}\frac{a_{k}}{2k+1/2}=\frac{\pi}{2y_{3}}.$ (4.2)
We observe that, even if it can occur that $a_{1}\in\Bigl(\frac{5\pi}{4y_{3}},\frac{3\pi}{2y_{3}}\Bigr)$, it still happens that, definitely in $k$, $a_{k}\in\Bigl(\frac{k\pi}{y_{3}},\frac{(2k+1/2)\pi}{2y_{3}}\Bigr)$. This is true because, recalling Eq.˜3.10, we have that
@@
and so there exists a $\overline{k}$, such that for all $k\geq\overline{k}$, $g<h$ when both evaluated at the middle point of the $k$-th interval. This fact together with Eq.˜3.11 implies that the unique solution is in the first half of the interval.
@@
Both sequences are vanishing, but $\{h(w_{l})\}$ does it faster than $\{g(w_{l})\}$, which implies that for sufficiently large $l$, $g(w_{l})>h(w_{l})$ and so the solution in the $(\overline{k}+l)$-th interval must be in $\Bigl(w_{l},\frac{(2(\overline{k}+l)+1/2)\pi}{2y_{3}}\Bigr)$. The arbitrarity of $w$ in $\Bigl(\frac{\overline{k}\pi}{y_{3}},\frac{(2\overline{k}+1/2)\pi}{2y_{3}}\Bigr)$ leads to Eq.˜4.2.
So we have that $a_{k}=\frac{(2k+1/2)\pi}{2y_{3}}+o(1)$ for $k\to+\infty$. (the symbol $o(1)$ here denotes a generic term vanishing as $k\to+\infty$. We use this property to estimate $\mathcal{N}_{-\Delta_{\alpha,y}^{\textup{D}}}(R)$ as $R\to+\infty$ excluding at most finitely many resonances. This is equivalent to find the highest $k$ for which
@@
$s=\sqrt{4y_{3}^{2}R^{2}-\ln^{2}{(s)}}+o(1).$ (4.3)
We see that $s<2y_{3}R$ and $s=O(R)$. This implies that $\ln^{2}{(s)}=O\Bigl(\ln^{2}(R)\Bigr)$. Now we substitute this in Eq.˜4.3, obtaining
@@
And so for $k$ holds ($\lfloor\cdot\rfloor$ and $\{\cdot\}$ indicate respectively the integer part and the fractional part) $k=\Biggl\lfloor\frac{y_{3}}{\pi}R-\frac{1}{4}\Biggr\rfloor.$
@@
$\mathcal{N}_{-\Delta_{\alpha,y}^{\textup{D}}}(R)=2\Biggl\lfloor\frac{y_{3}}{\pi}R-\frac{1}{4}\Biggr\rfloor+o(1)$ (4.4)
which implies equation (4.1) in the statement ∎
@@
We give a comparison with analogous results for standard Schrödinger operators $\Delta+V$. Firstly, for real potentials in $C^{\infty}_{\text{comp}}(\mathbb{R}^{3})$ (see [41]) or $L^{\infty}_{\text{comp}}(\mathbb{R}^{3})\cap H^{\frac{n-3}{2}}$ (see [51]) it is known that infinitely many resonances exist. Concerning asymptotic distribution of resonances, in [20] (see also [19]) it is shown that for Schrödinger operators $\Delta+V$ with generic $C^{\infty}_{\text{comp}}(\mathbb{R}^{n})$ one has
$\limsup_{R\to+\infty}\frac{\ln{\Bigl(\mathcal{N}_{-\Delta+V}(R)\Bigr)}}{\ln{R}}=n$
from which it follows the upper bound $\mathcal{N}_{-\Delta+V}\leq CR^{n}$, where the optimal bound $\mathcal{N}_{-\Delta+V}=CR^{n}$ is attained for radial potentials (see [54]), but not in general. The asymptotic behavior was already noticed in [40] for finitely many point interactions in $\mathbb{R}^{3}$, and compared with the analogous behavior on quantum graphs. See also the earlier work [53]. While the asymptotic distribution given by (4.1) seems somewhat anomalous in a three dimensional model, one has to take into account that point interactions can be recovered as limits of standard Schrödinger operators with rescaled smooth potentials, that however have to satisfy some very special conditions (see for example [3]).